OpenStudy (anonymous):

find the tangents to the curve y=2x^3 +3x at the points where the slope is 9. what is the smallest slope on the curve? at what value of x does the curve have this slope? please show all steps!

OpenStudy (anonymous):

differentiate the function to get the slope of the function at all points y' = 6x^2 +3 so 9 = 6x^2 +3 so x = +-1 if x is 1 put it into the original y = 2x^3 +3x y = 2 * 1^3 +3*1 y = 2 +3 y = 5 point is (1,5) sub -1 y = 2 * (-1)^3 +3(-1) y = -2 -3 y = -5 so point is (-1,5)

OpenStudy (anonymous):

second part. y' (the slope) is at a min i.e 0 when x = 0 so ans 0 and 0 !!!